I can get the anser to this but I dont know how they got it.v and r are riding their bikes.V rides at 10 miles per hour while R rides at 7.r has a 30 minute head start. How long has R been riding when V is 7 miles ahead of him?
Answer by edjones(7311)

d=ts
distance=time*speed
for R t=t s=7 d=7t
for V t=t-.5 d=d+7 s=10
R's distance=V's distance-7
7t=10(t-.5)-7
7t=10t-5-7
7t=10t-12 Then subtract 10t from each side
-3t=-12 then divide both sides by -3
t=4hrs Time R has been riding when V is 7 mi ahead of him
Check: 4*7=28mi=distance of R
====3.5*10=35mi=distance of V
EdJones